### v1

# Identifying and Mitigating API Misuse in Large Language Models

Abstract
API misuse in code generated by large language models (LLMs) represents a serious emerging challenge in software development. While LLMs have demonstrated impressive code generation capabilities, their interactions with complex library APIs remain highly prone to errors, potentially leading to software failures and security vulnerabilities. This paper presents the first comprehensive study of API misuse patterns in LLM-generated code, analyzing both method selection and parameter usage across Python and Java. Through extensive manual annotation of 3,892 method-level and 2,560 parameter-level misuses, we develop a novel taxonomy of four distinct API misuse types specific to LLMs, which significantly differ from traditional human-centric misuse patterns. Our evaluation of two widely used LLMs, StarCoder-7B (open-source) and Copilot (closed-source), reveals significant challenges in API usage, particularly in areas of hallucination and intent misalignment. We propose Dr.Fix, a novel LLM-based automatic program repair approach for API misuse based on the aforementioned taxonomy. Our method substantially improves repair accuracy for real-world API misuse, demonstrated by increases of up to 38.4 points in BLEU scores and 40 percentage points in exact match rates across different models and programming languages. This work provides crucial insights into the limitations of current LLMs in API usage and presents an effective solution for the automated repair of API misuse in LLM-generated code.
1 Introduction
Recent advances in large language models (LLMs) have significantly enhanced the sophistication of automatic code generation, as these models are trained on vast datasets of source code [1, 2, 3, 4]. However, according to previous work, LLMs may still generate suboptimal or incorrect API usage in code generation despite being exposed to extensive code during training [5]. Library APIs have been pivotal in software development, enabling complex functionalities across various domains, including network communication, data visualization, and scientific computation [6]. Moreover, as libraries evolve by introducing new features, deprecating old ones, or changing usage patterns, LLMs may struggle to adapt to these changes [7, 8]. With over 590,000 external libraries [^1] https://pypi.org/ in languages like Python, using the right API correctly can be a considerable challenge, even for experienced developers.
Despite these advancements, existing studies on API misuse in LLM-generated code have focused primarily on controlled, synthetic scenarios. For example, [9] investigated the quality of code generated by several LLMs using StackOverflow-style programming questions, finding that GPT-4 generated code with API misuses in 62% of cases when tested on 18 widely used Java APIs. Similarly, [10] studied API misuse in LLMs by examining security-related programming questions in Java, revealing that LLMs misused APIs in approximately half of the tasks. A more recent error analysis of 1,000 samples from existing benchmarks identified three primary error types, “AttributeError”, “TypeError”, and “ValueError”, commonly associated with API misuse in LLM-generated code. While these studies have demonstrated that specific LLMs can generate code with API misuse on limited libraries, APIs, and programming languages, their findings may not generalize to real-world programming contexts due to their reliance on human-curated questions and scenarios.
A more common scenario in real-world software development involves LLMs integrated within IDE environments, which assist developers in continuing or completing existing code [11]. LLMs are often prompted with an incomplete code snippet and asked to suggest an appropriate API call. In this context, LLMs face distinct challenges: They may erroneously hallucinate functions from other libraries, suggest irrelevant or incorrect APIs, or specify invalid parameters, even when familiar with the target library. These context-dependent completion tasks present challenges different from those of generating standalone code snippets in response to specific queries. While previous research has studied API misuse of LLMs on task-oriented synthetic datasets [9, 10], there remains a critical gap in understanding whether LLMs exhibit similar patterns of API misuse in real-world code completion. Specifically, it remains unclear whether LLMs make the same types of mistakes in API usage as human developers do. Moreover, investigating API misuse patterns across both open-source and closed-source models is essential, as open-source models often lag behind their closed-source counterparts in terms of performance. Current approaches to mitigating API misuse primarily focus on automatic detection using predefined misuse types and API documentation. However, these methods face significant limitations when applied to LLM-based code generation, which involves millions of APIs and varying contexts. Even when misuse is detected correctly, rule- and program-analysis-based automatic program repair (APR) techniques often fail to fix misused APIs due to their limited semantic understanding [12, 13]. This suggests that LLM-based APR approaches, which can leverage the semantic understanding of code, may be more suitable for addressing such issues.
This work investigates how LLMs misuse APIs and examines their behavior in real-world code-generation scenarios. Specifically, we evaluate two LLMs: StarCoder-7B [3], an open-source LLM for code completion, and Copilot, a fine-tuned closed-source model based on GPT-4 [2]. Both models are widely used in IDE environments such as VSCode [^2] https://code.visualstudio.com/. While previous human-centric API misuse research often studies patches from fixed commits [12, 13], it is not feasible to treat LLM-generated code in the same way, as LLMs generate outputs in various formats. We sample code snippets from GitHub in Python and Java to systematically assess how these models misuse APIs. We conduct program analysis to localize the API positions and provide the code context before these APIs as input to the models. Unlike existing studies, we assess the correctness of API invocation in two key areas: method selection and parameter usage. The first aspect evaluates whether LLMs can suggest an appropriate API name based on the code context, while the second examines whether LLMs can predict reasonable API parameters. These aspects correspond to typical use cases in IDE environments’ code filling [14] and left-to-right code completion [15].
To the best of our knowledge, we are the first to analyze complex API misuse in LLM-based code generation through extensive manual annotation of thousands of code snippets. Our systematic study reveals that LLMs struggle significantly with API misuse, particularly hallucination and intent misalignment, with distinct patterns occurring between Python and Java. While Copilot outperforms StarCoder, their shared error patterns indicate common challenges in API usage. To address these issues, we propose a novel LLM-based APR method that significantly improves repair accuracy. BLEU scores increase by up to 38.4 points, and exact match rates improve by up to 40 percentage points across different models and languages. The key contributions of this research include:
  • We propose the first taxonomy of API misuse in the era of LLMs, which covers four types, different from human-centric misuse. We identify (1) Intent misuse, where models select syntactically correct but functionally inappropriate APIs; (2) Hallucination misuse, where models generate non-existent API methods or parameters; (3) Missing item misuse, where models omit required API parameters or methods, and (4) Redundancy misuse, where models introduce unnecessary API calls or method chaining that may lead to inefficiencies or errors.
  • We manually annotated the LLM-generated outputs and provided a dataset that contains 3892 method-level misuses (1,884 in Java and 2,048 in Python) and 2560 parameter-level misuses (1,279 in Java and 1,281 in Python). For each misuse sample, we provide the category based on the taxonomy.
  • We design a new LLM-based approach, Detect-reason-Fix (Dr.Fix) to automatically repair the misuse cases introduced by LLMs, demonstrating its efficacy among strong LLM-based APR methods.
The rest of the paper is organized as follows. We outline the background in Section 2. We elaborate on the details of the experimental setting for the case study and introduce the misuse taxonomy in Section 3. Based on the annotations on generating two LLMs on Python and Java, we conduct quantitative analysis and share several insights in Section 4. Then, we propose the Dr.Fix for the in-wild API misuse repair and evaluate its performance against LLM-based APR on our benchmark in Section 5.
import torch
import torch.nn as nn
class GRUCell(nn.Module):
A simple implementation of a GRUCell in PyTorch.
def __init__(self, input_size, hidden_size, bias=True):
super(GRUCell, self).__init__()
self.input_size = input_size
self.hidden_size = hidden_size
self.bias = bias
# LLM is expected to complete the next line with an API call, such as nn.Linear
self.ih =
Figure 1: An example of LLM code completion suggesting a PyTorch API call for the last line.
2 Background
2.1 Motivation
Consider the example of LLM code generation shown in Figure 1. The LLM is prompted with an incomplete code snippet and asked to suggest an appropriate API call. The LLM typically makes several errors when suggesting code completion or insertion: the LLM generates incomplete method calls and/or with erroneous parameters; the LLM generates method calls to similar but unreleated API; the LLM generates extraneous method calls not needed in the target usage context; the LLM does not follow a required sequence of method calls to properly set up and use the target library due to its API’s complexity; and the LLM needs to invoke API calls to multiple libraries but fails to correctly use these different library APIs together. All of these errors can be very challenging to detect and fix, statically and dynamically.
2.2 API Misuse
APIs are essential in modern software systems, enabling developers to interact with third-party libraries and services. Correct API usage ensures that the functionality of a library is used as intended, while misuse can lead to errors, unexpected behavior, inefficiencies, or security vulnerabilities [16]. API misuse occurs when developers violate the intended usage constraints of an API, such as input types, function call sequences, or data handling practices. Unlike general programming errors (e.g., type mismatches, null pointer dereferencing) [17], which are language-specific, API misuse may not immediately result in failure but can cause subtle issues like incorrect outputs or degraded performance [18].
The causes of API misuse are multifaceted, including unclear documentation, lack of domain-specific knowledge, and breaking changes in evolving APIs [19]. The consequences can range from minor performance bottlenecks to severe failures, especially in high-risk domains like healthcare or finance [10]. Addressing API misuse requires better documentation, improved tooling, and clear guidelines to ensure developers follow best practices. Research typically focuses on external misuse of libraries in open-source codebases, often identified by manually inspecting bug-fixing commits [20, 21, 22].
In this study, we focus on the misuse of third-party library APIs arising from improper usage patterns or violations of API contracts rather than issues related to the underlying programming language or general software evolution. We adopt definitions from [23], with minor adaptations. An API is the public interface exposing software elements (e.g., classes and methods) to the outside world. An API element includes any constituent part of an API used in client code, such as methods and parameters. An API misuse is the violation of one or more API usage constraints, leading to misbehavior such as errors, crashes, incorrect output, or vulnerabilities. Proper understanding and adherence to API constraints are crucial for preventing these issues and ensuring the API functions as intended.
2.3 Mitigation of API Misuse
API usage can be categorized into two primary types: directly calling API methods or instantiating objects from API classes [24]. API misuses are defined as violations of the implicit or explicit usage constraints of APIs [16]. Recently, numerous API-misuse detectors have been developed [25, 26, 27]. Generally, there are three types of API misuse detection approaches: static detectors, which identify API misuses through static analysis of source or binary code [28]; dynamic detectors, which detect API misuses through dynamic analysis [29]; and hybrid approaches that combine mining techniques with static analysis [30]. Regardless of the detection technique, existing methods (1) require API specifications, considering the violation of these specifications as API misuse. However, these specifications are often incomplete and difficult to obtain [27]; or (2) do not require explicit API specifications, instead relying on the assumption that the majority usage pattern in a large-scale code corpus represents valid usage. An API is considered misused if it deviates from this majority pattern. The limitation is that this assumption may not always hold, especially for rarely used APIs where the majority pattern does not necessarily reflect valid usage [26]. Recently, there has been promising work exploring the use of LLMs to repair API misuse by fine-tuning paired commit data [13]. This approach outperforms previous automatic program repair methods, providing a more effective solution to mitigate API misuse in code generation.
3 Taxonomy and Qualitative Studies
This section introduces a new taxonomy for API misuse in LLMs, based on our qualitative studies. We begin by establishing the framework for code generation and then exploring various types of API misuse under the proposed taxonomy.
3.1 Research Questions
We wanted to answer the following key research questions about API misuse by current LLMs:
RQ1: How well do LLMs handle API invocation in code generation? We first wanted to investigate how well LLMs can correctly invoke APIs in code generation tasks, using widely-used automatic evaluation metrics.
RQ2: How do LLMs misuse APIs in the wild? As automatic evaluation metrics can only reflect the overall correctness, we then wanted to understand the specific misuse patterns of LLMs from the human perspective.
3.2 Evaluation Elements
To invoke an API correctly, LLMs are expected to predict both the correct API method name and pass valid arguments for the API parameters. We design the following two scenarios to evaluate the correctness of the generated API elements:
API Method Inference
In this evaluation setup, we mask out the API method name in a code snippet and ask the models to infer the appropriate method, similar to type inference [31, 32]. This evaluation assesses whether the models can comprehend the intent of the code context and recommend a suitable method name based on the imported packages and the parameters’ context.
import requests
def fetch_data(url):
# model will fill in the mask by predicting an API name
response = requests.[MASK](url)
return response.json()
Figure 2: Example for API Method Inference: The model should predict ‘get‘ as the appropriate method name.
As shown in Figure 2, the model is expected to predict ‘get‘ as the appropriate method name.
API Parameter Generation
In this case, rather than focusing on predicting the method name, we evaluate the models’ ability to generate valid and contextually appropriate parameter values for a given API method. Here, the model is provided with a partially completed code snippet where the API method name is already specified, but all the context (except for the opening parenthesis) is removed. The models are expected to complete the arguments correctly by recalling the parameter names and types associated with the API method and identifying potential variables as valid arguments.
import requests
def fetch_data():
url = "http://example.com/api/data"
response = requests.get([MASK])
return response.json()
Figure 3: Example for API Parameter Generation: The model should predict ‘url’ as the appropriate parameter.
As shown in Figure 3, the model is expected to predict ‘url‘ as the appropriate parameter.
3.3 LLM Models
LLMs typically differ in their architecture, size, and pretraining tasks. In practice, decoder-only models are preferred in the IDE environment, where developers utilize models to complete code and generate comments [33]. They may also commonly use the LLM to summarise code to aid understanding and to generate test cases for target code. Current IDEs like VSCode and Cursor [^3] https://www.cursor.com/ provide two types of decoder-only models for users: Open-source LLMs and Closed-source LLMs. Open-source LLMs are typically deployed in the user’s environment to optimize the inference speed (e.g., low latency) and flexibility (e.g., domain-specific fine-tuning) [34], while closed-source LLMs are more powerful but can only be inferred via company-hosted remote APIs. To study and compare representative examples of these different LLM models, we choose the widely used open-source model, StarCoder, and Copilot, a closed-source model provided by GitHub.
StarCoder is the first fully permissive LLM family trained on code by Hugging Face, outperforming popular LLMs and matching or surpassing closed-source models like ‘code-cushman-001‘ from OpenAI (the original Codex model that powered early versions of GitHub Copilot). With a context length of over 8,000 tokens, the StarCoder models can process more input than any other open-source LLM, enabling a wide range of interesting applications. In addition, the model supports both left-to-right code completion and fill-in-the-middle code prediction. Unlike previous LLMs that are only trained on common languages, StarCoder supports 86 programming languages from high resources (e.g., Java) to low resources (e.g., Dart). Furthermore, StarCoder provides three different model sizes, from 1B to 15B. The 1B model maximizes the inference speed, and the 15B model optimizes the generation quality. In our study, we chose StarCoder-7B, which is the best model balancing between accuracy and efficiency.
Copilot is a closed-source code suggestion model based on OpenAI’s GPT-4, initially launched in 2022. While the model is further fine-tuned for coding tasks, it has a much better understanding of natural language context than open-source LLMs like StarCoder. More than 46% of developers’ code files, on average, are reportedly generated by GitHub Copilot. Similar to StarCoder, Copilot offers two generation paradigms: code completion and code infilling. Besides, Copilot utilizes a client-side model to reduce the frequency of unwanted suggestions when they might prove disruptive to a developer’s workflow, helping it better respond to each developer using it. In this work, we use the Copilot version, which started in 2024, and manually paste code snippets into VSCode for model inference.
3.4 Hyperparameters
To align real-world software development in the IDE with LLMs, we use the default hyperparameters specified in each VSCode extension. Regarding StarCoder, we follow the configurations of Hugging Face’s llm-ls [^4] https://github.com/huggingface/llm-ls, a Language Server Protocol server leveraging LLMs for code completion. Specifically, we use the temperature of 0.2, the maximum new token number of 150, and the top_p value of 0.95. For GitHub’s Copilot model, we use the official v1.252.0 extension in VSCode 1.95.3 and prompt the model with the code snippet in a single file. Specifically, we directly query the models for autocompletion without any customized instructions.
3.5 Dataset
We describe how we construct the dataset for our qualitative study. We leverage The Stack [35] data, the largest and most up-to-date permissively licensed GitHub corpus containing source code files in 358 programming languages from GitHub repositories. Due to the high data quality resulting from the comprehensive filtering mechanisms, The Stack dataset has been partially utilized by a series of LLMs for pretraining. Specifically, we use v1.2, the latest version that opted out of the requested and malicious data, and exclude the subset used to train StarCoder to avoid data contamination issues [36].
To conduct a thorough investigation, we selected two popular programming languages, Python and Java. As each code file can contain multiple library APIs, we localize all the positions of APIs (methods and parameters) via static analysis. It is important to note that there could be alternative signatures for the same APIs, particularly where methods share the same name but differ in their parameters (e.g., overloading or different argument types). Additionally, we will consider APIs that have the same function or call methods but are located in different paths or namespaces due to different versions. To account for this, we ensure that all distinct signatures are included in our dataset. For example, an API could appear in various forms depending on the number and types of parameters used. To ensure comprehensive coverage of different programming scenarios, we include only one sample for each fully qualified-named API signature in our dataset, which helps us capture the diversity of API usage across various contexts. Furthermore, there are many duplicated APIs invoked in various code snippets. To ensure that our examination can cover diverse programming scenarios in the wild, we choose only one sample for each unique, fully-qualified, named API signature for each file. In total, we include 3,000 samples for both Python and Java, respectively, based on the most frequently used APIs. For each sample of API usage, we treat the code before the API method as a prefix and the parenthesized parameters as a suffix for Parameter Prediction and the code before the API parameter as a prefix for Parameter Generation, which results in 24,000 generated samples.
3.6 Evaluation Metrics
As we randomly sample our dataset from a raw GitHub code corpus without environment configurations (e.g., dependencies) and oracles, it is hard to assess the code quality via compilation and test cases. Following the Prior works [37, 38, 39], we use BLEU [40], CodeBERTScore [41], and Exact Match (EM) as the evaluation metrics to determine the (un)likelihood that the models can invoke the same APIs as human developers.
BLEU
BLEU calculates the percentage of 4-gram overlap between the reference fixed code and the candidate code generated by the models. For simplicity, we refer to it as BLEU in our work. The candidate fixed code represents the code generated by the models, while the reference fixed code represents the ground-truth code written by developers. BLEU is defined as follows:
$\text{BLEU}=\text{BP}\cdot\exp\left(\sum_{n=1}^{4}w_{n}\log p_{n}\right)$
where $p_{n}$ is the precision of n-grams, $w_{n}$ is the weight for n-grams (usually equal), and BP is the brevity penalty to account for shorter candidate sequences.
CodeBERTScore
CodeBERTScore, inspired by BERTScore [42], measures semantic consistency between generated and reference code by computing the cosine similarity of their token representations. Specifically, CodeBERTScore computes the F scores by combining the precision and recall. Precision ($\text{CodeBERTScore}_{P}$) is calculated as the average maximum similarity between each token in the generated code and tokens in the reference code, while recall ($\text{CodeBERTScore}_{R}$) is computed as the average maximum similarity between each token in the reference code and tokens in the generated code. These are combined using a weighted harmonic mean to produce the final score. In this work, we adopt $\text{CodeBERTScore}_{F_{3}}$, which prioritizes recall over precision, as it better aligns with functional correctness in code generation tasks. The formula is defined as:
$\text{CodeBERTScore}_{F_{3}}$ $=$
$\frac{10\cdot\text{CodeBERTScore}_{P}\cdot\text{CodeBERTScore}_{R}}{9\cdot\text{CodeBERTScore}_{P}+\text{CodeBERTScore}_{R}}$
Exact Match (EM)
EM is defined as the count of generated code instances $\hat{Y}$ whose abstract syntax trees (ASTs) exactly match the ASTs of the corresponding reference/ground-truth code $Y^{*}$. It is a strict metric that evaluates syntactic equivalence and may exclude reasonable outputs that differ in syntax but are semantically correct. EM serves as a lower bound, as different programs can achieve the same functionality but be written differently. In this context, $\hat{Y}$ represents the set of generated code instances, and $Y^{*}$ represents the set of reference or ground-truth code instances existing in The Stack dataset, which well captures how humans invoke those APIs. Each $\hat{y}$ is a single generated code instance, and each $y^{*}$ is a corresponding reference code instance. The EM metric counts the number of AST-based exact matches in the test set instead of the exact values passed into each API. The formula for EM is defined as:
$\texttt{EM}(\hat{Y},Y^{*})\triangleq\sum_{\hat{y}\in\hat{Y}}\texttt{matches}(\texttt{AST}(\hat{y}),\texttt{AST}(y^{*}))$
3.7 EM Failure Qualitative Example Analysis
We examined samples that failed on the EM score, indicating that they are highly likely to be misused by LLMs. We employed an iterative qualitative analysis approach inspired by grounded theory and qualitative content analysis. We randomly sampled 300 pairs of failed cases from each model, categorized by element type (method or parameter) and programming language (Python or Java), yielding a total of 1,200 failure samples. This sample size ensures statistical significance, achieving a 99% confidence level with a 5% margin of error. [^5] https://www.surveysystem.com/sscalc.htm
Two authors, each with over five years of programming experience in Python and Java, reviewed and categorized the API misuse samples. We apply open coding to label the errors descriptively without predefined categories. Through constant comparison, they iteratively grouped similar misuses into higher-level conceptual categories.
First, the two authors independently familiarized themselves with the dataset by reviewing a subset of API misuse cases. Each annotator labeled 50 randomly selected samples, took detailed notes on their observations, During this process, they were allowed to search online for API information when necessary to validate the correctness of the method and parameter usage. Following this initial review, the annotators conducted axial coding, systematically grouping similar misuses into broader conceptual categories based on shared characteristics. They then met to compare their annotations, resolve discrepancies, and iteratively refine their categorization framework. Through multiple review cycles, they refined their labels, clarified ambiguous definitions, and introduced new categories where necessary. In the end, they progressively consolidated similar misuse patterns into four distinct categories: Intent Misuse, Hallucination, Missing Parameter, Redundant API Call. With the finalized codebook, the two authors independently labeled the remaining API misuse cases. To assess inter-annotator agreement, we computed Cohen’s Kappa score [43], achieving a final value of 0.97, indicating almost perfect agreement. The annotation process spanned 200 hours, encompassing multiple iterations of refinement to ensure accuracy and consistency.
It is important to note that while we focus on four main categories of API misuse, there are other potential misuse types, such as security-related or deprecation issues. However, these types are more knowledge-intensive and cannot be easily identified without sufficient contextual information or detailed code snippets. For instance, detecting deprecated methods or security vulnerabilities requires understanding the specific API versions and security considerations, which cannot be reliably determined from code snippets alone. Therefore, our analysis focuses on categories that are more directly identifiable through the provided code and are less dependent on external, domain-specific knowledge. We present each category, which is introduced below, with illustrative examples.
import numpy as np
from blmath.numerics import vx
def farthest(from_point, to_points):
Find the farthest point among the inputs, to the given point.
Return a tuple: farthest_point, index_of_farthest_point.
- absolute_distances = np.abs(to_points - from_point)
+ absolute_distances = vx.magnitude (to_points - from_point)
index_of_farthest_point = np.argmax(absolute_distances)
farthest_point = to_points[index_of_farthest_point]
return farthest_point, index_of_farthest_point
Figure 4: A diff-formatted example of intent misuse: Incorrect use of np.abs instead of vx.magnitude in Python.
3.8 API Misuse Classification
Type 1: Intent Misuse
Intent misuse occurs when a model selects an API or method that is syntactically correct but does not align with the intended functionality or task. This leads to incorrect behavior due to the model’s misunderstanding of the method’s purpose. In Figure 4, the method farthest() is intended to find the farthest point from a reference point. To calculate this, the absolute distances between points must be computed. The model incorrectly uses np.abs, a method from the numpy library, which computes the element-wise absolute values of an array. Although syntactically correct, np.abs does not calculate the magnitude of the vectors, which is the intended purpose in this case. The correct API is vx.magnitude, which is designed to compute the magnitude of vectors and is thus more suitable for the task. Therefore, the misuse here stems from the model’s failure to select the appropriate method that aligns with the intended operation, resulting in a mismatch between the code’s functionality and its purpose.
import java.util.Collections;
public static List<String> getMenuTitles(final Resources res) {
final List<String> menuList = new ArrayList<String>();
- strArray = res.getStringArray(R.array.navigation_main_menu_titles);
+ strArray = res.getStringArray(R.array.navigation_main_titles);
Collections.addAll(menuList, strArray)
if (!FeatureConfig.USE_OPENDATA) {
menuList.remove(MENU_STATISTICS_INDEX);
return menuList;
import us.ihmc.commons.MathTools;
@Override
public void initialize() {
currentTime.set(0.0);
- parameterPolynomial.setCubic(trajectoryTime.getDoubleValue(), 0.0, Double.POSITIVE_INFINITY);
+ MathTools.checkIntervalContains(trajectoryTime.getDoubleValue(), 0.0, Double.POSITIVE_INFINITY);
parameterPolynomial.setQuintic(0.0, trajectoryTime.getDoubleValue(), 0.0, 0.0, 0.0, 1.0, 0.0, 0.0);
currentOrientation.set(initialOrientation);
currentAngularVelocity.setToZero();
currentAngularAcceleration.setToZero();
Figure 5: Diff-formatted Examples of hallucination misuse: navigation_main_menu_titles instead of navigation_main_titles and non-existent setCubic instead of checkIntervalContains in Java.
Type 2: Hallucination Misuse
Hallucination occurs when a model introduces an entirely incorrect API method or parameter that does not exist or is not required for the task. The model may create or "hallucinate" an API call that seems plausible but invalid in the context, leading to erroneous behavior or a compilation/runtime error. Concurrently, several studies have studied hallucination issues of LLM-based code generation [44, 45, 46]. However, they consider any bugs made by LLMs to be a type of hallucination. In this work, we define “Hallucination” as the case when LLMs generate non-existent API parameters or methods in the wild. In Figure 5, the model incorrectly uses the method res.getStringArray(R.array.
navigation_main_menu_titles)
as a parameter for the Collections.addAll() method. The correct call should be res.getStringArray(R.array.
navigation_main_titles)
, where the parameter name was likely fabricated by the model, leading to an API misuse. This is a case of parameter hallucination, where the model introduces a nonexistent or incorrect parameter in the API call. In Figure 5, the model wrongly selects parameterPolynomial.setCubic(trajectoryTime.
getDoubleValue(), 0.0, Double.POSITIVE_INFINITY)
when the correct method should have been
MathTools.checkIntervalContains() for checking a time interval. Here, the model replaces the correct API method with a non-existent or irrelevant one, resulting in method hallucination. The model mistakenly introduces an incorrect method that does not align with the context or expected API behavior, leading to a significant deviation from the intended functionality.
import org.chromium.base.PathUtils;
public static void loadLibrary() {
Context appContext = ContextUtils.getApplicationContext();
- PathUtils.setPrivateDataDirectorySuffix(PRIVATE_DATA_DIRECTORY_SUFFIX);
+ PathUtils.setPrivateDataDirectorySuffix(PRIVATE_DATA_DIRECTORY_SUFFIX, appContext);
try {
LibraryLoader libraryLoader = LibraryLoader.get(LibraryProcessType.PROCESS_WEBVIEW);
libraryLoader.loadNow(appContext);
libraryLoader.switchCommandLineForWebView();
} catch (ProcessInitException e) {
throw new RuntimeException("Cannot load WebView", e);
Figure 6: A diff-formatted example of missing item misuse: Omission of the required appContext parameter in setPrivateDataDirectorySuffix in Java.
Type 3: Missing Item Misuse
Missing item misuse occurs when the model omits a required API method or parameter, which results in incomplete or incorrect behavior. The model may fail to include a necessary parameter or method, leading to errors or malfunctioning of the system. This type of misuse typically occurs when the model does not fully understand the API’s requirements or ignores important details that are essential for correct usage. In Figure 6, the model incorrectly predicts the API usage by omitting a necessary parameter from the method PathUtils.setPrivateDataDirectorySuffix(). The expected method call should include both the PRIVATE_DATA_DIRECTORY_SUFFIX constant and the appContext parameter. The model fails to include appContext, which is required for the method to properly set the private data directory suffix. This omission leads to parameter omission misuse, where the model overlooks a crucial parameter in the API call, causing potential runtime errors or unexpected behavior.
import org.chromium.base.PathUtils;
public static void loadLibrary() {
Context appContext = ContextUtils.getApplicationContext();
LibraryLoader libraryLoader = LibraryLoader.get(LibraryProcessType.PROCESS_WEBVIEW);
try {
// Redundant method chaining
- libraryLoader.loadNow(appContext).initialize(); // initialize() is unnecessary
+ libraryLoader.loadNow(appContext); // Correct: loadNow() is sufficient
libraryLoader.switchCommandLineForWebView();
} catch (ProcessInitException e) {
throw new RuntimeException("Cannot load WebView", e);
Figure 7: A diff-formatted example of redundancy misuse: Unnecessary chaining of initialize() after loadNow() in Java. The correct implementation should avoid redundant method calls where a single method is sufficient.
Type 4: Redundancy
Redundancy misuse occurs when the model includes unnecessary method chaining or repeated API calls, which do not contribute to the task and lead to inefficiency or potential errors. In the example provided in Figure 7, the model incorrectly chains the initialize() method after loadNow(), even though loadNow() already performs the necessary initialization. This redundant method of chaining can lead to confusion and potential side effects. The correct behavior would involve using only the necessary method call, loadNow(), which is sufficient for the task.
4 Results
4.1 RQ1: How well do LLMs handle API invocation in code generation?
TABLE I: Performance comparison of StarCoder and Copilot in Python and Java for method and parameter prediction tasks.
Model Python Java
BLEU CodeBERTScore EM BLEU CodeBERTScore EM
Method StarCoder 84.6 92.3 2,268 81.2 86.4 1,328
Copilot 88.3 92.6 1,911 86.1 87.7 1,593
Parameters StarCoder 85.0 92.8 2,312 82.0 87.0 1,330
Copilot 89.0 93.0 1,956 87.0 88.2 1,617
We first wanted to investigate how well our two studied LLMs can generate code to correctly invoke APIs in code generation. As described above, we sample 3,000 distinct code snippets in Python and Java, respectively, from The Stack dataset. After prompting StarCoder and Copilot to complete Parameter Prediction and Parameter Generation independently, we collect 24,000 samples in total. We measure the correctness of the API invocation using three metrics introduced in subsection 3.6: BLUE, CodeBERTScre, and EM. Table I shows the performance of StarCoder and Copilot on in-the-wild API usage tasks in Python and Java. The results reveal that Copilot consistently outperforms StarCoder across all metrics, with the EM metric providing the clearest indication of model capability in generating syntactically correct and precise code.
For Python, Copilot achieves an EM of 1911, compared to StarCoder’s 2268. This significant difference underscores Copilot’s superior ability to produce exact matches to reference implementations. Additionally, Copilot surpasses StarCoder in BLEU and CodeBERTScore, achieving scores of 88.3 and 92.6, respectively, compared to StarCoder’s 84.6 BLEU and 92.3 CodeBERTScore. These results highlight that Copilot is more effective in generating both syntactically accurate and semantically similar Python code. The higher EM score further solidifies its advantage in tasks requiring exact code generation.
In the Java domain, the trend remains consistent, with Copilot outperforming StarCoder in all metrics. For Java, Copilot achieves an EM of 1593, compared to StarCoder’s 1328, again reflecting its ability to generate more precise API usage code. Similarly, Copilot’s BLEU score of 86.1 and CodeBERTScore of 87.7 surpass StarCoder’s BLEU of 81.2 and CodeBERTScore of 86.4. These results suggest that Copilot is better suited for Java tasks, producing fluent and semantically aligned code with higher precision.
The comparison between these two decoder-based models highlights the significance of pretraining strategies and dataset scale. Copilot’s consistently higher EM scores across both Python and Java suggest it is better optimized for tasks that demand exact code matches, while its performance in BLEU and CodeBERTScore indicates strong fluency and semantic alignment. Although StarCoder performs competitively in CodeBERTScore, its lower EM scores in both languages suggest room for improvement in generating syntactically exact outputs.
437StarCoderCopilot1721235
(a) Wrong Method Inference in Java
190StarCoderCopilot547542
(b) Wrong Method Inference in Python
665StarCoderCopilot3781005
(c) Wrong Parameter Generation in Java
237StarCoderCopilot593451
(d) Wrong Parameter Generation in Python
Figure 8: Comparison of incorrect API usage (wrong method inference and parameter generation) between StarCoder and Copilot in Java and Python.
Figure 8 shows the overlap and discrepancies in API misuse between StarCoder and Copilot. These Venn diagrams illustrate instances where one model produces incorrect API usage, either in the method call or parameter context, for Python and Java. In Python, Copilot demonstrates a clear advantage over StarCoder in API method inference, as shown in 8b. StarCoder has 190 unique errors, compared to Copilot’s 547, with 542 errors shared between the two models. These shared errors highlight common challenges, likely stemming from ambiguous requirements or limitations in the training datasets. (8a) shows a similar trend for Java, where StarCoder produces 437 unique method errors versus Copilot’s 172, with 1235 shared errors. This suggests that both models face consistent difficulties in method inference due to the complexity of Java APIs. For parameter generation, the patterns differ between languages. In Python (8d), StarCoder has 237 unique parameter errors, while Copilot has 593, with 451 shared errors. In Java (8c), StarCoder generates 665 unique parameter errors compared to Copilot’s 378, with 1005 shared errors, indicating distinct challenges in precise parameter usage across both languages.
{findingbox}
Finding 1: While Copilot demonstrates higher BLEU and CodeBERTScore metrics across both languages, error patterns vary significantly by language. Copilot shows fewer API misuses than StarCoder in Java but makes more errors in Python for both method inference and parameter generation. Both models share a substantial number of common errors, indicating persistent challenges in handling complex API usage scenarios.
4.2 RQ2: How do LLMs misuse APIs in the wild?
TABLE II: Comparison of API Misuse by StarCoder in Java and Python. “None” misuse type means that the LLM-based API usage is considered appropriate based on human inspection.
Misuse Type Java Python
Params (%) Methods (%) Params (%) Methods (%)
Intent 5.5 27.0 7.4 30.4
Hallucination 43.3 34.6 31.3 40.9
Redundancy 8.2 10.2 14.8 5.3
Missing Item 15.0 11.3 26.5 3.0
None 28.0 16.9 20.0 20.4
TABLE III: Comparison of API Misuse by Copilot in Java and Python. “None” misuse type means that the LLM-based API usage is considered appropriate based on human inspection.
Misuse Type Java Python
Params (%) Methods (%) Params (%) Methods (%)
Intent 4.0 22.0 6.0 25.0
Hallucination 39.0 29.0 25.0 33.0
Redundancy 7.0 12.0 12.0 7.0
Missing Item 15.0 12.5 29.0 7.7
Acceptable 35.0 24.5 28.0 27.3
We further analyze the patterns and distributions of API misuse of LLMs. Specically, we study the annotated outputs from both StarCoder and Copilot in subsection 3.7, which failed on EM score. Table II and Table III present a detailed analysis of how different types of API misuse manifest in Java and Python code generated by StarCoder and Copilot, respectively. The tables categorize the misuses into several types and compare the percentage occurrences of each type for both programming languages.
In both Java and Python, the most common misuse type is Hallucination, with significant differences in frequency between the two languages. For StarCoder, hallucinations account for 43.3% of parameter misuses and 34.6% of method misuses in Java, while for Python, the hallucination rate is even higher: 31.3% for parameters and 40.9% for methods. For Copilot, hallucinations are slightly lower but still prominent, with 39.0% of parameter misuses and 29.0% of method misuses in Java, and 25.0% for parameters and 33.0% for methods in Python. This suggests that while both models experience hallucination-related errors, Copilot generally performs better, especially in reducing method misprediction errors.
Figure 9: Workflow of Dr.Fix to repair the API misuse.
Intent misuses, which occur when the model generates technically correct APIs but fails to fulfill the intended task, also appear frequently. In Python, this misuse type appears in 30.4% of method predictions and 7.4% of parameter predictions for StarCoder. For Copilot, these numbers are lower, at 25.0% for methods and 6.0% for parameters. In Java, intent misuse is present in 27.0% of method misuses and 5.5% of parameter misuses for StarCoder, compared to 22.0% for methods and 4.0% for parameters for Copilot. This indicates that while both models sometimes misunderstand the intended usage of APIs, Copilot shows better performance in both languages.
The Redundancy misuse, where extra or unnecessary parameters or methods are generated, is another notable error. In Python, redundancy misuses account for 14.8% of parameter errors and 5.3% of method errors for StarCoder, compared to 12.0% for parameters and 7.0% for methods for Copilot. In Java, redundant mistakes occur at a rate of 8.2% for parameters and 10.2% for methods for StarCoder, while Copilot has lower rates of 7.0% for parameters and 12.0% for methods. Redundancy errors typically arise when the model includes superfluous API arguments or calls, sometimes due to overfitting on training data patterns or confusion over which parameters are necessary for a given method.
The Missing Item misuse, where expected parameters or method calls are omitted entirely, is a significant issue in Python for both models, appearing in 26.5% of parameter-related errors for StarCoder and 29.0% for Copilot. In Java, missing parameter misuse is less common, with 15.0% for StarCoder and 15.0% for Copilot. This suggests that Python may pose additional challenges to the models when it comes to identifying all required inputs, possibly due to its dynamic typing system or the variation in how functions and methods are structured.
{findingbox}
Finding 2: In general, the most frequent API misuse in LLM-generated code is hallucinations. Hallucination and intent misuses are more prominent in method generation, with Python experiencing higher rates of hallucination errors. Copilot generally performs better than StarCoder, with higher acceptable rates and lower intent and hallucination misuses. Redundancy and missing item misuses are common in both languages, especially for parameters, with Python showing a unique challenge in the form of missing parameter-related misuses.
5 Mitigation
Mitigating LLM-based API misuse in real-world applications presents several challenges that must be effectively addressed. While automatic misuse detection methods exist, they often necessitate significant manual effort to rectify API misuse. This underscores the need for an automatic approach to aid developers in repairing API misuse. However, the realm of APR methods tailored specifically for API misuse remains largely unexplored. The only relevant work is [13], where the authors evaluated various learning-aid APR models by fine-tuning pre-trained language models on a large-scale (human-written) git commit dataset. However, the study only empirically studied APR for Java and cannot be extended to various programming languages, including Python.
Repairing many LLM-based API misuses requires a nuanced semantic understanding, particularly in deciphering the implicit intents within the code context. Traditional APR methods, reliant on language-specific static analysis or heuristic rules, struggle without delving into these contextual nuances. Moreover, the nondeterministic and cross-lingual nature of libraries utilized in source code poses a formidable challenge for knowledge-constrained APR methods, rendering them ineffective in addressing API misuse across diverse programming environments.
We investigate a promising avenue for API misuse repair - Prompt-based LLM-based APR. Unlike conventional learning-based APR methods trained on task-specific datasets, Prompt-based LLM-based APR is poised to offer superior generalization across a myriad of prompts. Previous research has shown that LLMs can self-repair/debug, enhancing the quality of generated content [47].
5.1 Dr.Fix: Repairing API Misuse via Detection and Reasoning
Simply prompting LLMs to repair API misuses may not always yield reliable results. LLM-based APR methods can struggle with understanding the nuanced context of the code and the specific intent behind API usage. These models may generate repairs that are syntactically correct but semantically inappropriate, failing to align with the developer’s original intentions. Moreover, LLMs may lack the capability to consistently identify and address the underlying issues causing the misuse. To address these limitations, we propose a novel approach: Mitigating API Misuse via Detect-reason-Fix (Dr.Fix).
Figure 9 overviews our Dr.Fix approach, which integrates detection and reasoning capabilities through multiple stages of LLM prompting. The process begins with the detection stage, where LLMs are prompted to recognize potential API misuse within the code. The LLM is prompted with "Identify any incorrect or unsafe API usages in the following code snippet." Once potential misuses are detected, the taxonomy classification stage follows. Here, the LLM classifies these misuses into predefined categories, such as intent and hallucination misuse. In the reasoning stage, the LLM is prompted to validate the detected misuse with a detailed analysis, considering the broader context and the code’s intended functionality. This process is motivated by self-verification reasoning [48], which has been demonstrated in natural language processing. Finally, in the repair suggestion stage, the LLM generates repair suggestions based on its analysis. These suggestions are crafted to be syntactically correct and semantically meaningful, aligning with the developer’s original intentions.
Incorporating Dr.Fix into the development workflow can significantly reduce the manual effort required to rectify API misuses. By automating both the detection and repair processes, Dr.Fix enhances the reliability and efficiency of software development. This novel approach represents a significant advancement in automated program repair, particularly for LLM-based API misuse, and underscores the potential of combining LLM capabilities with sophisticated reasoning techniques to address complex programming challenges.
5.2 Backbone Instruction-tuned Models
Llama-3.1
Llama 3.1 [49] is a collection of multilingual LLMs, a collection of pre-trained and instruction-tuned generative models in 8B, 70B, and 405B sizes released by Meta AI. The Llama 3.1 instruction-tuned text-only models (8B, 70B, 405B) are optimized for multilingual dialogue use cases and outperform many available open-source and closed-chat models on common industry benchmarks. Compared to the original Llama 3 series, Llama-3.1 is significantly improved in terms of math and reasoning capabilities. In our evaluation, we use Llama-3.1 70B, a variant achieving state-of-the-art (SOTA) performance on various benchmarks.
Qwen2.5-Coder
Qwen2.5-Coder [50] is a series of code-specific generative models designed to excel in a wide range of programming tasks such as code generation, completion, reasoning, and repair. Built upon the Qwen2.5 architecture, these models come in sizes ranging from 0.5B to 32B parameters and have been pre-trained on a vast corpus of over 5.5 trillion tokens. Through a combination of extensive data cleaning, scalable synthetic data generation, and balanced data mixing, Qwen2.5-Coder achieves SOTA performance across more than 10 code-related benchmarks. It consistently outperforms other models of similar size, showcasing impressive capabilities in code intelligence while retaining strong general and math skills. With its permissive licensing, Qwen2.5-Coder is positioned to drive innovation in code-related AI research and be widely adopted by developers in real-world applications. We chose Qwen2.5-Coder 32B, the most capable one among the series.
GPT-4o
GPT-4o [^6] https://openai.com/index/hello-gpt-4o/ is a highly advanced variant of OpenAI’s GPT-4 model, fine-tuned specifically for enhanced efficiency and specialized performance across various natural language processing and multimodal tasks. Available in different configurations, GPT-4o delivers superior reasoning, problem-solving, and contextual understanding capabilities compared to its predecessors. With a focus on optimizing performance while reducing computational overhead, GPT-4o excels in both complex language tasks and more specialized use cases, such as coding, translation, and content generation. Leveraging a blend of advanced training techniques, including reinforcement learning from human feedback, GPT-4o sets a new standard for AI models in terms of accuracy, responsiveness, and versatility, making it ideal for real-world applications across industries.
5.3 Hyperparameters
In this study, we set a maximum token length to 2048, a temperature of 0, and a top-p of 0.95. The few-shot learning approach [51] is used to prompt the models across four stages: detection, classification, reasoning, and repair suggestion. Specifically, we use two shots per stage as a demonstration to ensure the output format. For each type inside the API misuse taxonomy, we sample two failure cases to teach the models.
The prompts for each stage are designed as follows: (1) Detect — “Identify any incorrect usages in the following code snippet based on one of the following categories: Intent, Hallucination, Redundancy, Missing.” The definition of each category is provided after the instructions. To prevent LLMs from hallucinating API misuse, we append “Please answer with "No Issue" if there is no obvious API misuse.” at the end of the prompt. (2) Reason — “Think step by step and explain why this API usage is incorrect and how it deviates from the intended usage.” (3) Fix — “Propose a correction that aligns with the intended functionality of the API.”
The models’ performance is evaluated based on BLEU, CodeBERTScore, EM accuracy, and refusal rate to assess their ability to detect and repair API misuses across different categories and programming languages.
The refusal rate is introduced as an additional evaluation metric inspired by [52, 53, 54]. It measures the model’s ability to avoid hallucinating an API misuse when none exists in the code snippet. This score is defined as the percentage of instances where the model correctly refuses to detect a misuse when no misuse is present in the code. The formula for the refusal rate is:
Refusal Rate=Number of Correct RejectionsTotal Number of No Misuse $\text{Refusal Rate}=\frac{\text{Number of Correct Rejections}}{\text{Total Number of No Misuse Samples}}\times 100$
TABLE IV: 10% Sample of Misuse Types for Java and Python
Misuse Type Java Python
Params # Methods # Params # Methods #
Intent 11 55 10 39
Hallucination 89 71 40 53
Redundancy 17 21 19 7
Missing 31 23 34 4
Total 149 133 105 98
5.4 Dataset
We sampled 10% of code snippets from each annotated misuse type in Section 3. As shown in Table IV, the dataset includes a representative sample of API misuse types across Java and Python. Specifically, the dataset captures the following misuse types: Intent, Hallucination, Redundancy, Position, and Missing. For each type, a proportional number of samples was extracted based on their original distribution. To further enhance the evaluation of repair approaches, including Dr.Fix and other baselines, we introduced an additional 100 snippets that exactly match the reference source code. This addition allows us to test the ability of repair tools to avoid unnecessary hallucinations or misinterpretations of correct API usage. By including these reference-aligned snippets, we can measure the precision of tools in distinguishing between actual misuse and correct implementations.
5.5 Results Analysis
TABLE V: Performance of Models across Different Methods. Dr.Fix (w/o taxonomy) means that the models are not prompted with the defined API misuse taxonomy and demonstrated examples, and perform the detection based on their own understanding. Baseline denotes the LLM-based APR method where the models are prompted to generate the fixed code snippet if they detect the API misuse inside the given snippet.
Model Method BLEU CodeBERTScore EM
Java – Method
Llama-3.1 Baseline 30.5 25.7 35
Dr.Fix (w/o taxonomy) 65.2 66.3 75
Dr.Fix 68.9 69.2 80
Qwen-Coder-2.5 Baseline 36.3 28.4 24
Dr.Fix (w/o taxonomy) 55.1 60.2 92
Dr.Fix 58.3 62.6 96
GPT-4o Baseline 57.8 46.1 52
Dr.Fix (w/o taxonomy) 76.1 65.3 92
Dr.Fix 79.4 68.2 96
Java – Parameter
Llama-3.1 Baseline 53.6 64.6 34
Dr.Fix (w/o taxonomy) 72.3 85.1 72
Dr.Fix 75.8 88.3 78
Qwen-Coder-2.5 Baseline 55.7 73.3 21
Dr.Fix (w/o taxonomy) 73.2 89.5 80
Dr.Fix 76.4 92.3 85
GPT-4o Baseline 62.7 82.6 26
Dr.Fix (w/o taxonomy) 80.1 89.2 84
Dr.Fix 82.5 91.5 88
Python – Method
Llama-3.1 Baseline 34.6 58.4 25
Dr.Fix (w/o taxonomy) 54.2 87.6 55
Dr.Fix 57.8 90.3 60
Qwen-Coder-2.5 Baseline 15.1 43.8 14
Dr.Fix (w/o taxonomy) 49.3 81.2 48
Dr.Fix 52.6 84.6 52
GPT-4o Baseline 47.0 69.5 31
Dr.Fix (w/o taxonomy) 60.5 90.1 66
Dr.Fix 63.7 92.7 71
Python – Parameter
Llama-3.1 Baseline 78.7 75.7 12
Dr.Fix (w/o taxonomy) 82.1 84.3 40
Dr.Fix 84.5 86.8 45
Qwen-Coder-2.5 Baseline 81.2 62.3 10
Dr.Fix (w/o taxonomy) 85.1 87.2 52
Dr.Fix 87.4 89.4 56
GPT-4o Baseline 84.8 67.2 11
Dr.Fix (w/o taxonomy) 86.9 90.8 34
Dr.Fix 88.6 92.4 38
The performance of the models across different methods is summarized in Table V. The results highlight the significant improvement in API misuse detection and repair capabilities achieved by incorporating the Dr.Fix methodology, with and without the taxonomy. This section provides an in-depth analysis of the findings for both Java and Python across the evaluated metrics.
Java Performance: For Java, both Dr.Fix and Dr.Fix (w/o taxonomy) outperform the baseline method across all metrics. In the Method misuse category, GPT-4o with Dr.Fix achieves the highest BLEU score, CodeBERTScore, and EM, demonstrating its ability to generate syntactically and semantically appropriate repairs. The intermediate version, Dr.Fix (w/o taxonomy), also shows substantial improvements, with GPT-4o achieving a BLEU score of 76.1 and an EM of 92, which further enhance to 79.4 and 96, respectively, with the full taxonomy. Similarly, Qwen2.5-Coder and Llama-3.1 show marked improvements with both versions of Dr.Fix, with BLEU scores increasing from 36.3 to 58.3 and from 30.5 to 68.9, respectively. In the Parameter misuse category, the results follow a similar trend. GPT-4o achieves the best performance with BLEU, CodeBERTScore, and EM. Dr.Fix (w/o taxonomy) notably enhances the performance of all models, with improvements in both syntactic and semantic alignment, as reflected in the BLEU and CodeBERTScore metrics, which are further boosted by the full taxonomy.
Python Performance: For Python, both Dr.Fix and Dr.Fix (w/o taxonomy) demonstrate significant improvements over the baseline method. In the Method misuse category, GPT-4o achieves the highest BLEU score and CodeBERTScore, indicating superior generalization and repair accuracy. The EM score also improves from 31 to 71 with the full taxonomy, while Dr.Fix (w/o taxonomy) achieves an EM of 66. Similarly, Qwen2.5-Coder and Llama-3.1 achieve notable gains in BLEU and CodeBERTScore metrics with both versions of Dr.Fix, showcasing the effectiveness of the methodology in improving code repair suggestions. In the Parameter misuse category, all models exhibit enhanced performance with Dr.Fix. GPT-4o achieves a BLEU score of 88.6 and a CodeBERTScore of 92.4 with the full taxonomy, surpassing other models in both accuracy and precision. While the baseline method performs reasonably well in this category, both versions of Dr.Fix provide a clear advantage, ensuring that repair suggestions align more closely with developer intentions, with the full taxonomy offering the most significant improvements.
Figure 10: Refusal rate comparison of Llama-3.1, Qwen-Coder-2.5, and GPT-4o with and without Dr.Fix.
Refusal Rate: The refusal rate metric evaluates the models’ ability to avoid hallucinating API misuses when no misuse exists. As shown in Figure 10, Dr.Fix significantly enhances the refusal rates across all tested models. GPT-4o achieved the highest refusal rate of 84% with Dr.Fix, demonstrating its exceptional ability to differentiate between correct and incorrect API usage. Similarly, Qwen-Coder-2.5 and Llama-3.1 showed substantial improvements, achieving refusal rates of 67% and 46%, respectively, compared to their baseline configurations, which had refusal rates of 0%, 55%, and 35%, respectively. Dr.Fix (w/o taxonomy) also showed promising results, with refusal rates of 70%, 55%, and 35% for GPT-4o, Qwen-Coder-2.5, and Llama-3.1, respectively. Notably, Dr.Fix slightly decreases the refusal rate when no API misuse taxonomy is given for the model, suggesting that it effectively reduces false positives by avoiding unnecessary refusals for correct API usage. These results highlight Dr.Fix’s effectiveness in reducing hallucinated API misuses across diverse LLM architectures, with the full taxonomy version providing the most significant improvements. {findingbox} Our experimental results demonstrate that Dr.Fix significantly enhances API misuse repair across all models, achieving up to 130% improvement in exact match rates compared to baseline approaches while maintaining high code quality. The framework reduces refusal errors by 14-20% absolute percentage points across different architectures, showcasing its effectiveness as a reliable program repair solution for diverse programming languages.
6 Threats to Validity
6.1 Internal Validity
One potential threat to internal validity is the reliance on subjective human annotations for the initial classification of API misuse. Although two annotators with over five years of programming experience independently reviewed the samples, biases in understanding or interpretation of the API misuse may have influenced the categorization. We mitigated this by ensuring high inter-annotator agreement, as indicated by the high Cohen’s Kappa score, and resolving disagreements through detailed discussions. Another threat arises from the inherent variability in LLM outputs due to stochastic sampling. To address this, we kept the model parameters fixed across experiments, varying only the prompts as necessary. However, given resource constraints, experiments were conducted once, and results may vary under repeated evaluations. Future studies should consider multiple iterations to assess consistency and variability in repair suggestions.
6.2 External Validity
A significant threat to external validity is the generalizability of the findings to other programming languages, APIs, and LLMs. Our dataset is limited to Python and Java, focusing on specific API misuse categories. While these languages and categories are widely used and relevant, the results may not directly extend to other contexts, such as low-level languages or highly specialized APIs. Similarly, the LLMs evaluated in this study, including GPT-4o, Qwen2.5-Coder, and Llama-3.1, represent a subset of available models. The effectiveness of Dr.Fix with other LLMs or on alternative programming datasets remains uncertain. Future research should expand the scope to include additional languages, APIs, and models to validate the framework’s broader applicability.
6.3 Construct Validity
Construct validity may be threatened by the choice of evaluation metrics and dataset design. Metrics such as BLEU and CodeBERTScore, while widely used, primarily measure syntactic and semantic similarity and may not fully capture the functional correctness of the repairs. LLMs-as-Judges [55, 56] is hence a more promising direction to explore for code quality evaluation. Moreover, the inclusion of 100 snippets without API misuse to evaluate refusal rates introduces a potential bias in the dataset composition. Although this addition aims to measure the models’ ability to avoid hallucinating misuse, its proportion relative to actual misuse cases may influence the results. Future work could explore alternative metrics or evaluation methods that better assess functional correctness and developer usability.
7 Related Work
7.1 Large Language Models for Code
LLMs have emerged as a transformative approach for code generation tasks, broadly categorized into standard language models and instruction-tuned models. Standard language models are pre-trained on raw corpora using next-token prediction, exemplified by models such as Codex [57], CodeGen [58], CodeGeeX [59], CodeT5 [60] and StarCoder [3]. Codex, with 12B parameters, and StarCoder, featuring 15.5B parameters, demonstrated significant advancements in program synthesis through training on GitHub code and additional sources. Instruction-tuned models represent the next evolution, with examples like ChatGPT employing Reinforcement Learning with Human Feedback [61]. Open-source alternatives include WizardCoder [62], which fine-tunes StarCoder using techniques like Evol-Instruct, and InstructCodeT5+ [63], which extends CodeT5 through instruction-tuned processes. These developments have enabled zero-shot capabilities in code generation tasks, demonstrating emergent behaviors that were previously unattainable while maintaining the essential citations and academic format.
7.2 Bug Study of Large-Language-Model-Generated Code
Understanding bugs in LLM-generated code is crucial for improving code generation capabilities. While prior studies have analyzed bugs in deep learning systems broadly [64, 65, 66], recent research has focused specifically on defects in LLM-generated code [67]. [68] investigated bugs in non-standalone code generated by LLMs, though their study was limited by potential data leakage issues and the use of now-outdated models. [69] addressed these limitations by evaluating LLM-generated code with a robust benchmark designed to minimize data leakage, providing a more accurate representation of model performance in practical applications. Our work builds upon these insights by focusing specifically on API misuse patterns in LLM-generated code, introducing a taxonomy of error types and proposing the Dr.Fix repair mechanism.
7.3 Program Repair by Large Language Models
The success of LLMs in Natural Language Processing has led to their application in Automated Program Repair (APR) [70, 71]. Key advancements include integrating LLMs with additional contextual information, such as TFix [72], which builds on the T5 model [73] by incorporating error messages from diagnostic tools to improve repair accuracy. Specialized models for code review and repair have also emerged [74, 75], along with zero-shot learning approaches that frame repair as a cloze-style task. Notable examples include Xia and Zhang’s work [71] using masked language modeling and CIRCLE [76]’s prompt-based approach for code completion. Our work, Dr.Fix, builds upon these foundations by specifically addressing API misuse in generated code, combining detection, reasoning, and repair capabilities across Python and Java codebases.
8 Conclusion and Future Work
In this work, we systematically studied API misuse in LLM-generated code, analyzing thousands of code snippets across Python and Java. Through rigorous manual annotation, we developed a comprehensive taxonomy of API misuse types and quantified their prevalence in popular code generation models. Our findings reveal that LLMs struggle significantly with hallucination and intent misalignment, with distinct patterns emerging between programming languages. The shared error patterns between Copilot and StarCoder highlight fundamental challenges in API usage that persist across different model architectures. We introduced Dr.Fix, a novel LLM-based APR approach that combines detection and reasoning capabilities to address these challenges. Our evaluation demonstrates that Dr.Fix significantly outperforms existing learning-based APR methods, substantially improving repair accuracy and maintaining high refusal rates for correct API usage. The results underscore the effectiveness of our staged approach in understanding and fixing API misuse.
Future work should focus on extending Dr.Fix’s capabilities to handle more complex API misuse scenarios and investigating its applicability to other programming languages. Additionally, exploring the integration of Dr.Fix with development environments and investigating ways to improve its reasoning capabilities through enhanced prompting strategies could further advance automated API repair. We also plan to expand our taxonomy and dataset as new patterns of API misuse emerge with the evolution of code generation models.